OpenStudy (anonymous):

can i get some help, im pretty sure i know how to do this problem but can someone check it?

OpenStudy (anonymous):

whats your question?

OpenStudy (anonymous):

x/5 2/3 5/y

OpenStudy (anonymous):

are these the answer mads?

OpenStudy (anonymous):

tht was the question after that i get 3x+2y=25

OpenStudy (anonymous):

hmmm well r u just supposed to fill in for x? or is it like combine like terms?

OpenStudy (anonymous):

i have to find x and y

OpenStudy (anonymous):

alright gimme a min to figure it out

OpenStudy (anonymous):

alright your answer is 5

OpenStudy (anonymous):

3(5) + 2(5) = 25

OpenStudy (anonymous):

for x or y?

OpenStudy (anonymous):

so x and y are both 5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it gives the answer of 25 and im not sure you could fit another number it ones place and get the same answer.

OpenStudy (anonymous):

but then it wuld be 5/5=2/3=5/5.... tht doesnt work O.o

OpenStudy (anonymous):

x/3=2/3=5/y

OpenStudy (anonymous):

x and y

OpenStudy (anonymous):

You must set Y equal to zero to find X and vise-versa. So To find X, your question will look like this: 3x+2(0)=25 And to solve for Y: 3(0)+2y=25

OpenStudy (anonymous):

Mads was given that answer already, anothonys. It isn't correct..

OpenStudy (anonymous):

i got x=8.3and y-12.5

OpenStudy (anonymous):

dude, back down

OpenStudy (anonymous):

now...

OpenStudy (anonymous):

stop being a cutie pie ;)

OpenStudy (anonymous):

back down and be a man!

OpenStudy (anonymous):

doomdude ;) you lookin hot ;) <3

OpenStudy (anonymous):

you will be reported if you do not!

OpenStudy (anonymous):

.... Ever hear of chivalry?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

thats not very nice :(

OpenStudy (anonymous):

well i see my question has turned unproductive.....

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i will try to help

OpenStudy (anonymous):

ya someone get a moderator in here

OpenStudy (anonymous):

....

OpenStudy (anonymous):

too late to remove your posts anthony...

OpenStudy (anonymous):

i still don't understand my own questin

OpenStudy (anonymous):

um, i am sure amistre can help further

OpenStudy (anonymous):

he is good with explaining

OpenStudy (anonymous):

hey @mads4566 since you are only given one equation you have to solve this by trial and error

OpenStudy (amistre64):

ok, my first question is; what was the original question?

OpenStudy (anonymous):

3x+2y=25 there are infinite many solutions how many solutions are you looking for?

OpenStudy (anonymous):

look at that first question as two different questions: X 2 5 2 --- = --- AND --- = --- 5 3 Y 3

OpenStudy (anonymous):

x/5=2/3=5/y is the originl

OpenStudy (anonymous):

i have to find x and y

OpenStudy (amistre64):

so its a group of proportions is what it looks like to me

OpenStudy (amistre64):

cross multiplying will seem like a good bet

OpenStudy (anonymous):

OO so x=10/3 and y = 15/2

OpenStudy (anonymous):

use that equation maker thingy :)

OpenStudy (anonymous):

how could x possibly be 10/3?

OpenStudy (anonymous):

yeah @amistre64

OpenStudy (amistre64):

x/5=2/3 ; multiply by 3*5 to clear the fractions 3x = 10; divide off the 3 x = 3/10 2/3=5/y ; multiply thru by 3*y to clear the fractions 2y = 15; divide off teh 2 y = 15/2

OpenStudy (anonymous):

so i need to sollve for one at a time instead or extending cross multiplication ... like x/5=2/3 and solve for x

OpenStudy (anonymous):

your x is wrong bro

OpenStudy (amistre64):

its simpler to do them one at a time

OpenStudy (anonymous):

Oohhh

OpenStudy (amistre64):

yeah, typoed it :) x = 10/3

OpenStudy (anonymous):

Thank you so much!! :)

OpenStudy (amistre64):

youre welcome, the idea is to make them all equal, and since you have a solid value to work with; equate the unknowns to that one

OpenStudy (amistre64):

finicky; x is any value that makes it work :)